OpenStudy (anonymous):

for which equation is x=-1/2 a solution? a. 7-4x=-9 b. 9x+2=7 c. 6+8x=-10 d. 10x-3=-8

OpenStudy (anonymous):

just substitute in all the and find out!

OpenStudy (anonymous):

i cant!

OpenStudy (anonymous):

-4x=-9+7 -4x=-2 4x=2 x=1/2

OpenStudy (anonymous):

why not?? what do you get when x= -1/2 for 7-4x?

OpenStudy (anonymous):

im slow :(

OpenStudy (anonymous):

why don't you try answering what i asked in my previous post?

OpenStudy (anonymous):

okay it's 7-4x=-9 right?

OpenStudy (anonymous):

can you help me with this one "at her job, mary works 4 hours on monday, 3 hours on wednesday, and 6 hours on saturday. her boss adds a $10 bonus to her paycheck, which then totals $120.50 . what is mary's hourly pay? a. $8.50 b.$9.27 c. $5.24 d. $8.05" IM REALLLY STUCK ON THIS ONE!